# Feature Fusion from Head to Tail: an Extreme Augmenting Strategy for Long-Tailed Visual Recognition

### v1

@@ 3.4 Rationale Analysis
We explore the theoretical rationality of the seemingly intuitive operation of H2T in depth in this section. For ease of analysis, we assume that the feature maps to be fused are rearranged in order without loss of generality. Then, after the pooling layer, the feature can be written as $f_{i}^{T}=[\dot{f}_{i}^{T},\ddot{f}_{i}^{T}]$, where $\dot{f}_{i}$ and $\ddot{f}_{i}$ denote the portions of features that are retained and to be fused, respectively $w_{i}^{T}=[\dot{w}_{i}^{T},\ddot{w}_{i}^{T}]$ is the corresponding classifier weights.
@@
However, as shown in Figure 1b and analyzed in Section 3.2, numerous tail class samples are incorrectly classified into head classes. Therefore, the trained model actually predicts $z_{h}>z_{t}$, namely,
@@
We use $\ddot{\theta}_{*}$ to represent the angle between $\ddot{w}_{*}$ and $\ddot{f}_{t}-\ddot{f}_{h}$. Eq. (7) is further simplified to
@@
Eq. (8) forces $w_{h}$ to move closer to tail class samples while pushing $w_{t}$ further away from them, thus increasing the distribution span of head classes. Conversely, Eq. (10) exerts an opposite force. Figure 3 geometrically interprets the rationale of H2T. Notably, when the fusion ratio $p$ is small, Eq. (10) exerts a greater force, enabling the classifier to be well-calibrated. As $p$ grows, the item $\ddot{w}_{t}^{T}\ddot{f}_{h}>\ddot{w}_{h}^{T}\ddot{f}_{h}$ in Eq. (7) progressively becomes more dominant. Under this case, even though Eq. (8) encourages $w_{t}$ to move away from tail class samples, which is larger than the force of Eq. (10), $\ddot{w}_{t}^{T}\ddot{f}_{h}>\ddot{w}_{h}^{T}\ddot{f}_{h}$ attracts $w_{t}$ towards head class samples, thus broadens the distribution span of tail classes. Consequently, the classifier increases its performance in the tail class, regardless of the value of $p$. Section 4.4 also experimentally validates this analysis.
Figure 3: Rationale analysis of H2T. $\bullet$: $x_{h}$. $\blacktriangle$: $x_{t}$. $\rightarrow$: $w_{h}$. $\rightarrow$: $w_{t}$. $\rightarrow$: $f_{t}-f_{h}$. Forces 1⃝ and 2⃝ are generated by Eq. (10) and Eq. (8), respectively. 1⃝ $>$ 2⃝ makes the tail class sample to“pull" closer to $w_{t}$ and “push" further away from $w_{h}$. This leads to the enlargement of the tail class space and a reduction in the priority of head class.
@@ 4 Experiments
Table 1: Comparison results on CIFAR10/100-LT w.r.t. top-1 accuracy (%).
Dataset CIFAR10-LT CIFAR100-LT
Backbone: ResNet-32 Imbalance Ratio
Method Year 200 100 50 200 100 50
@@
CE loss - 67.24 70.70 74.81 35.99 39.55 45.40
LDAM-DRW [4] 2019 73.52 77.03 81.03 38.91 42.04 47.62
DR+MU [19] 2020 73.06 79.15 84.21 41.73 45.68 50.86
MisLAS [56] 2021 77.17 82.12 85.55 43.45 46.71 52.31
${}^{\star}$ [31] 2022 - 81.00 86.60 - 45.80 52.60
GCL [27] 2022 78.56 82.06 85.36 44.76 48.61 53.55
${}^{\star}$ [36] 2022 - - - - 43.90 48.30
BSCE [${}^{\star}$ [36] 2022 - - - - 46.60 51.40
@@
MisLAS+H2T - 78.05 ($\uparrow$ 0.88) 83.23 ($\uparrow$ 1.11) 86.72 ($\uparrow$ 1.17) 43.84 ($\uparrow$ 0.39) 47.62 ($\uparrow$ 0.91) 52.73 ($\uparrow$ 0.42)
GCL+H2T - 79.42 ($\uparrow$ 0.86) 82.40 ($\uparrow$ 0.34) 85.41 ($\uparrow$ 0.05) 45.24 ($\uparrow$ 0.48) 48.88 ($\uparrow$ 0.27) 53.76 ($\uparrow$ 0.21)
@@
BBN [57] 2020 73.47 79.82 81.18 37.21 42.56 47.02
RIDE [47] 2021 - - - 45.84 50.37 54.99
ACE [${}^{\star}$ 2021 - 81.20 84.30 - 49.40 50.70
ResLT [10] 2022 75.29 80.58 84.07 44.26 48.73 53.81
@@
ResLT+H2T - 77.25 ($\uparrow$ 1.96) 81.77 ($\uparrow$ 1.19) 84.99 ($\uparrow$ 0.92) 46.18 ($\uparrow$ 1.92) 49.60 ($\uparrow$ 0.87) 54.39 ($\uparrow$ 0.58)
  • Note: The results are obtained by re-implementing the official codes or reproducing with the given parameters, except the methods marked by $\star$ that are quoted from the corresponding papers.
  • The best and the second-best results are shown in underline bold and bold, respectively.
4.1 Datasets and Metrics
Dataset: We evaluated our proposed method on five widely-used benchmarks: CIFAR10/100-LT [4], ImageNet-LT [33], iNaturalist 2018 [42], and Places-LT [33]. CIFAR10/100-LT are the small-scale datasets that are sub-sampled from their balanced versions, namely, CIFAR10/100 [21]. We used the imbalance ratios ($\rho=\frac{n_{max}}{n_{min}}$) 200, 100, and 50 [7]. The original versions of imageNet-2012 [40] and Places 365 [59] are also balanced datasets. We use the same settings as Liu et al. [33] to obtain the long-tailed versions. iNaturalist is a natural dataset collected from all over the world and is heavily imbalanced. The 2018 version [42] is utilized in our experiment.
Metrics: Besides top-1 classification accuracy, following Liu et al. [33], the accuracy on three partitions: head ($n_{i}>100$), medium ($20<n_{i}\leq 100$) and tail ($n_{i}\leq 20$) on large-scale datasets are also compared.
4.2 Basic Settings
SGD with a momentum of 0.9 is adopted for all datasets. Stage II is trained with 10 epochs. For CIFAR10/100-LT, we refer to the settings in Cao et al. [4] and Zhong et al. [56]. The backbone network is ResNet-32 [13]. Stage I trains for 200 epochs.The initial learning rate ($l_{r}$) is 0.1 and the batch size is 128. $l_{r}$ is decayed at the $160^{th}$ and $180^{th}$ epochs by 0.1. For imageNet-LT and iNaturalist 2018, we use the commonly used ResNet-50, while for Places-LT, we utilize the ResNet-152 pre-trained on imageNet. For other settings, we mainly follow the original papers of the comparison methods. We reproduce the prior methods that provide the hyper-parameters for a fair comparison. For those methods that do not provide hyperparameters or code, we directly report their results presented in their original paper.
4.3 Comparisons to Previous Methods
Comparison Methods: We compare the proposed H2T with previous state-of-the-art (SOTA) methods. Two-stage methods, i.e., LDAM-DRW [4], decoupling representation (DR) [19], MisLAS [55] and GCL [27] are compared. DR is the first to decouple the learning of representation and classifier. Zhang et al. [55] and Zhong et al. [56] found that MixUp (MU) [53] can enhance representation learning, therefore we reproduce decoupling representation cooperated with MU for comparison. We also compare with the latest proposed data augmentation methods, i.e., MBJ [31] and CMO [36]. CMO is also an additive approach. We report the results of CE loss and balanced softmax cross-entropy loss (BSCE) [38] with CMO (abbreviated as CE+CMO and BSCE+CMO, respectively). The results of H2T accompanied by DR (CE loss is utilized), MisLAS, and GCL are reported. For multi-expert models, we compare BBN [57], RIDE [47], ACE [3], and ResLT [10]. Both RIDE and ResLT use 3-expert architectures. For RIDE in imageNet-LT and iNaturalist 2018, we report the results of 90 and 180 epochs (eps for short). For Places-LT, we report the results of RIDE with linear classifier and cosine classifier. The results of H2T integrated with RIDE and ResLT are presented for comparison.
Comparison Results: The results on CIFAR10/100-LT are shown in Table 1. H2T can further improve both single and multi-expert models. Applying H2T to the vanilla training of DR with CE loss (DR+H2T) enhances the performance by a significant margin, surpassing most recent methods. The biggest improvement exceeds 5% (on CIRAR-10-LT with $\rho=200$). H2T can also further boost the performance of two-stage SOTA methods, i.e., MisLAS and GCL, albeit with less significant improvements than for DR. H2T also bolsters the efficacy of multi-expert models. For example, RIDE and ResLT integrating with H2T outperform the original methods by 1.01% and 0.87% on CIFAR100-LT with $\rho=100$.
@@ 4.3 Comparisons to Previous Methods
Backbone: ResNet-50
Method Year Head Med Tail All
@@
CE loss - 64.91 38.10 11.28 44.51
LDAM-DRW [4] 2019 58.63 48.95 30.37 49.96
DR+MU [19] 2020 62.93 49.77 33.26 52.18
MisLAS [56] 2021 62.53 49.82 34.74 52.29
${}^{\star}$ [31] 2022 61.60 48.40 39.00 52.10
GCL [27] 2022 62.24 48.62 52.12 54.51
${}^{\star}$ [36] 2022 67.00 42.30 20.50 49.10
BSCE [${}^{\star}$ [36] 2022 62.00 49.10 36.70 52.30
@@
BBN [57] 2020 - - - 48.30
RIDE [47] (90 eps) 2021 64.33 51.73 36.14 54.08
RIDE [47] (180 eps) 2021 69.59 53.06 30.09 55.72
ACE [${}^{\star}$ 2021 - - - 54.70
ResLT [10] 2022 59.39 50.97 41.29 52.66
@@
RIDE+H2T (180 eps) - 67.55 54.95 37.08 56.92 ($\uparrow$ 1.20)
ResLT+H2T - 62.29 52.29 35.31 53.39 ($\uparrow$ 0.73)
  • Note: The same as Table 1.
@@
Backbone: ResNet-50
Method Year Head Med Tail All
@@
CE loss - 76.10 69.05 62.44 66.86
LDAM-DRW [4] 2019 - - - 68.15
DR+MU [19] 2020 72.88 71.15 69.24 70.49
MisLAS [56] 2021 72.52 72.08 70.76 71.54
${}^{\star}$ [31] 2022 - - - 70.00
GCL [27] 2022 66.43 71.66 72.47 71.47
${}^{\star}$ [36] 2022 76.90 69.30 66.60 68.90
BSCE [${}^{\star}$ [36] 2022 68.80 70.00 72.30 70.90
@@
BBN [57] 2020 - - - 69.70
RIDE [47] (90 eps) 2021 72.05 73.34 70.23 71.83
RIDE [47] (180 eps) 2021 76.52 74.23 70.45 72.80
ACE [${}^{\star}$ 2021 - - - 72.90
ResLT [${}^{\star}$ 2022 64.85 70.64 72.11 70.69
@@
RIDE+H2T (180 eps) - 75.69 74.22 71.36 73.11 ($\uparrow$ 0.31)
ResLT with H2T - 68.41 72.31 72.09 71.88 ($\uparrow$ 1.19)
  • Note: The same as Table 1.
@@
Backbone: ResNet-152
Method Year Head Med Tail All
@@
CE loss - 46.48 25.66 8.09 29.43
DR+MU [19] 2020 41.66 37.79 32.77 37.40
MisLAS [56] 2021 41.95 41.88 34.65 40.38
${}^{\star}$ [31] 2022 39.50 38.20 35.50 38.10
GCL [27] 2022 38.64 42.59 38.44 40.30
@@
LFME+OLTR [48] ${}^{\star}$ 2020 39.30 39.60 24.20 36.20
RIDE [47] (LC) 2021 44.79 40.69 31.97 40.32
RIDE [47] (CC) 2021 44.38 40.59 32.99 40.35
ResLT [10] 2022 40.30 44.40 34.70 41.00
@@
  • Note: CC, cosine classifier. LC, linear classifier. Others are the same as Table 1.
@@
We further validate H2T on large-scale datasets, including imageNet-LT, iNaturalist 2018 and Places-LT. The results are summarized in Table 2, Table 3 and Table 4, respectively. The baseline (CE loss) achieves high accuracy on frequent classes while the performance on rare classes is not satisfactory. Compared with the baseline, all methods except CMO and RIDE achieve a significant improvement in the medium and tail classes but at the expense of reduced accuracy on head classes. CMO uses the background of head classes to augment tail classes without reducing the training samples in head classes. CMO accompanying basic CE loss, therefore, improves model performance for both head and tail classes. However, CMO+CE shows less competitive results in improving the classification performance of tail classes. H2T can deliver consistent performance improvements for both single and multi-expert models, particularly for tail and medium classes. Nevertheless, the accuracy improvements of cosine classifier (CC) are less pronounced than those of linear classifier (LC). For example, we can observe that H2T improves GCL that uses CC much less than DR and MisLAS that use LC. In addition, introducing the smooth class margin in GCL can alleviate the sample aliasing between classes in the embedding space. Thus, the enhancement effect of H2T on GCL is less obvious. For example, GCL+H2T improves only by 0.11 on imageNet-LT. Table 4 further compares the improvements of H2T on RIDE with LC (by 1.06%) and CC (0.95%). Note that the margin on LC is more profound.
4.4 Analysis
In this section, We visualize the augmentation effect of H2T on each class, and the experiments are conducted on CIFAR10-LT for a more convenient and clearer presentation. Additionally, we investigate the effects of the sampler used for the fusing branch and the fusion ratio $p$, which are conducted on CIFAR100-LT. All ablations are performed with $\rho$ set to 100, and the DR+H2T method is employed.
Visualization of Feature Space: Figure 4 shows the t-SNE visualization of the distribution in embedding space, which demonstrates our motivation in Section 3.2 that H2T can enrich tail classes so that prevents over-squeezing and overfitting. The features are extracted from class-balanced sampling data. We can see that the distribution of tail classes is sparser than that of head classes. H2T enriches the diversity of intra- and inter-class tail classes with the same number of samples without external information, effectively alleviating the problem of tail class over-squeezing. It is worth noting that, a clear margin, namely the left of Figure 4 is superior in a balanced dataset since the classifier is unbiased for each class. In an imbalanced dataset, a clear margin gives the classifier room to squeeze tail classes in order to correctly classify the head class samples with overwhelming samples. Filling the marge with semantic samples can prevent over-squeezing for tail classes.
Figure 4: Feature distribution comparison w.o. and w. H2T.
The Influence of Sampler for Fusing Branch:

### v2

@@ 3.4 Rationale Analysis
Although the proposed H2T appears intuitive and straightforward on the surface, it is built on a foundation of rationale. We explore its theoretical rationality in depth in this section. For ease of analysis, we assume that the feature maps to be fused are rearranged in order without loss of generality. Then, after the pooling layer, the feature can be written as $f_{i}^{T}=[\dot{f}_{i}^{T},\ddot{f}_{i}^{T}]$, where $\dot{f}_{i}$ and $\ddot{f}_{i}$ denote the portions of features that are retained and to be fused, respectively. $w_{i}^{T}=[\dot{w}_{i}^{T},\ddot{w}_{i}^{T}]$ is the corresponding classifier weights.
@@
However, as shown in Figure 3, numerous tail class samples are incorrectly classified into head classes. Therefore, the trained model actually predicts $z_{h}>z_{t}$, namely,
@@
Figure 3: Frequency of predictive labels for tail classes samples. A large number of tail class samples are incorrectly recognized as head classes.
Figure 4: Rationale analysis of H2T. Forces 1⃝ and 2⃝ are generated by Eq. (9) and Eq. (7), respectively. 1⃝ $>$ 2⃝ makes the tail sample to “pull” closer to $w_{t}$ and “push” further away from $w_{h}$, leading to the adjustment of decision boundary and enlargement of the tail class space.
@@
We use $\ddot{\theta}_{*}$ (*** is $h$ or $t$) to represent the angle between $\ddot{w}_{*}$ and $\ddot{f}_{t}-\ddot{f}_{h}$. Eq. (6) is further simplified to
@@
Eq. (7) forces $w_{h}$ to move closer to tail class samples while pushing $w_{t}$ further away from them, thus increasing the distribution span of head classes. Conversely, Eq. (9) exerts an opposite force. Figure 4 geometrically interprets the rationale of H2T. Notably, when the fusion ratio $p$ is small, Eq. (9) exerts a greater force, enabling the classifier to be well-calibrated. As $p$ grows, the item $\ddot{w}_{t}^{T}\ddot{f}_{h}>\ddot{w}_{h}^{T}\ddot{f}_{h}$ in Eq. (6) progressively becomes more dominant. Under this case, even though Eq. (7) encourages $w_{t}$ to move away from tail class samples, which is larger than the force of Eq. (9), $\ddot{w}_{t}^{T}\ddot{f}_{h}>\ddot{w}_{h}^{T}\ddot{f}_{h}$ attracts $w_{t}$ towards head class samples, thus broadens the distribution span of tail classes. Consequently, the classifier increases its performance in the tail class, regardless of the value of $p$. We also experimentally validate this analysis Section 4.4.
@@ 4 Experiments
4.1 Datasets and Metrics
We evaluate our proposed H2T on four widely-used benchmarks: CIFAR100-LT (Cao et al. 2019), ImageNet-LT (Liu et al. 2019), iNaturalist 2018 (Van Horn et al. 2018), and Places-LT (Liu et al. 2019). CIFAR100-LT is a small-scale dataset that is sub-sampled from the balanced version CIFAR100 (Krizhevsky, Hinton et al. 2009). We used the imbalance ratios ($\rho=\frac{n_{max}}{n_{min}}$) 200, 100, and 50 (Chu et al. 2020). The original versions of imageNet-2012 (Russakovsky et al. 2015) and Places 365 (Zhou et al. 2017) are also balanced datasets. We use the same settings as Liu et al. (Liu et al. 2019) to obtain the long-tailed versions. iNaturalist is collected from all over the world and is naturally heavily imbalanced. The 2018 version (Van Horn et al. 2018) is utilized in our experiment. We also report the comparison results on CIFAR10-LT (Cao et al. 2019) in the Appendix. As for the metrics, besides top-1 classification accuracy, following Liu et al. (Liu et al. 2019), the accuracy on three partitions: head ($n_{i}>100$), medium ($20<n_{i}\leq 100$) and tail ($n_{i}\leq 20$) on large-scale datasets are also compared.
4.2 Basic Settings
SGD with a momentum of 0.9 is adopted for all datasets. Stage II is trained with 10 epochs. For CIFAR100-LT, we refer to the settings in Cao et al. (Cao et al. 2019) and Zhong et al. (Zhong et al. 2021). The backbone network is ResNet-32 (He et al. 2016). Stage I trains for 200 epochs. The initial learning rate is 0.1 and is decayed at the $160^{th}$ and $180^{th}$ epochs by 0.1. The batch size is 128. For imageNet-LT and iNaturalist 2018, we use the commonly used ResNet-50. For Places-LT, we utilize the ResNet-152 pre-trained on imageNet. We reproduce the prior methods that provide the hyper-parameters for a fair comparison. For those methods that do not provide hyperparameters or official codes, we directly report their results presented in the original paper.
4.3 Comparisons to Existing Methods
Compared Methods: For single model, we compare the proposed H2T with the following three kinds of methods: two-stage methods, i.e., LDAM-DRW (Cao et al. 2019), decoupling representation (DR) (Kang et al. 2020) [^1] We reproduce DR with MixUp for a fair comparison., MisLAS (Zhang et al. 2021), and GCL (Li, Cheung, and Lu 2022); decision boundary adjustment method, i.e., Adaptive Bias Loss (ABL) (Jin, Lang, and Lei 2023); data augmentation methods, i.e., MBJ (Liu, Li, and Sun 2022), and CMO (Park et al. 2022). We report the results of CE loss and balanced softmax cross-entropy loss (BSCE) (Ren et al. 2020) with CMO (abbreviated as CE+CMO and BSCE+CMO, respectively). The results of H2T accompanied by DR (CE loss is utilized), MisLAS, and GCL are reported. For multi-expert models, we compare BBN (Zhou et al. 2020), RIDE (Wang et al. 2021b), ACE (Cai, Wang, and Hwang 2021), and ResLT (Cui et al. 2023). Both RIDE and ResLT use 3-expert architectures. For Places-LT, we report the results of RIDE with linear and cosine classifiers.
Table 1: Comparison results on CIFAR100-LT.
Imbalance Ratio 200 100 50
@@ 4.3 Comparisons to Existing Methods
CE loss 35.99 39.55 45.40
LDAM-DRW (Cao et al. 2019) 38.91 42.04 47.62
DR+MU (Kang et al. 2020) 41.73 45.68 50.86
MisLAS (Zhong et al. 2021) 43.45 46.71 52.31
${}^{\star}$ (Liu, Li, and Sun 2022) - 45.80 52.60
GCL (Li, Cheung, and Lu 2022) 44.76 48.61 53.55
${}^{\star}$ (Park et al. 2022) - 43.90 48.30
${}^{\star}$ (Park et al. 2022) - 46.60 51.40
${}^{\star}$ (Jin, Lang, and Lei 2023) - 46.80 52.10
@@
MisLAS+H2T 43.84 47.62 52.73
GCL+H2T 45.24 48.88 53.76
@@
BBN (Zhou et al. 2020) 37.21 42.56 47.02
RIDE (Wang et al. 2021b) 45.84 50.37 54.99
${}^{\star}$ (Cai, Wang, and Hwang 2021) - 49.40 50.70
${}^{\star}$ (Park et al. 2022) - 50.00 53.00
ResLT (Cui et al. 2023) 44.26 48.73 53.81
@@
ResLT+H2T 46.18 49.60 54.39
  • Note: The backbone is ResNet-32. $\star$ denotes the results quoted from the corresponding papers. The best and the second-best results are shown in underline bold and bold, respectively.
@@
Method Head Med Tail All
@@
CE loss 64.91 38.10 11.28 44.51
LDAM-DRW (Cao et al. 2019) 58.63 48.95 30.37 49.96
DR (Kang et al. 2020) 62.93 49.77 33.26 52.18
MisLAS (Zhong et al. 2021) 62.53 49.82 34.74 52.29
${}^{\star}$ (Liu, Li, and Sun 2022) 61.60 48.40 39.00 52.10
GCL (Li, Cheung, and Lu 2022) 62.24 48.62 52.12 54.51
${}^{\star}$ (Park et al. 2022) 67.00 42.30 20.50 49.10
${}^{\star}$ (Park et al. 2022) 62.00 49.10 36.70 52.30
${}^{\star}$ (Jin, Lang, and Lei 2023) 62.60 50.30 36.90 53.20
@@
BBN (Zhou et al. 2020) - - - 48.30
RIDE (Wang et al. 2021b) 69.59 53.06 30.09 55.72
${}^{\star}$  (Cai, Wang, and Hwang 2021) - - - 54.70
${}^{\star}$ (Park et al. 2022) 66.40 53.90 35.60 56.20
ResLT (Cui et al. 2023) 59.39 50.97 41.29 52.66
@@
ResLT+H2T 62.29 52.29 35.31 53.39
  • Note: The backbone is ResNet-50. Others are the same as Table 1.
@@
Method Head Med Tail All
@@
CE loss 76.10 69.05 62.44 66.86
LDAM-DRW (Cao et al. 2019) - - - 68.15
DR (Kang et al. 2020) 72.88 71.15 69.24 70.49
MisLAS (Zhong et al. 2021) 72.52 72.08 70.76 71.54
${}^{\star}$ (Liu, Li, and Sun 2022) - - - 70.00
GCL (Li, Cheung, and Lu 2022) 66.43 71.66 72.47 71.47
${}^{\star}$ (Park et al. 2022) 76.90 69.30 66.60 68.90
${}^{\star}$ (Park et al. 2022) 68.80 70.00 72.30 70.90
${}^{\star}$ (Jin, Lang, and Lei 2023) - - - 71.60
@@
BBN (Zhou et al. 2020) - - - 69.70
RIDE (Wang et al. 2021b) 76.52 74.23 70.45 72.80
ACE (Cai, Wang, and Hwang ${}^{\star}$ 72.90
RIDE+ ${}^{\star}$ (Park et al. 2022) 68.70 72.60 73.10 72.80
ResLT (Cui et al. ${}^{\star}$ 64.85 70.64 72.11 70.69
@@
ResLT with H2T 68.41 72.31 72.09 71.88
  • Note: The backbone is ResNet-50. Others are the same as Table 1.
@@
Method Head Med Tail All
@@
CE loss 46.48 25.66 8.09 29.43
DR (Kang et al. 2020) 41.66 37.79 32.77 37.40
MisLAS (Zhong et al. 2021) 41.95 41.88 34.65 40.38
${}^{\star}$ (Liu, Li, and Sun 2022) 39.50 38.20 35.50 38.10
GCL (Li, Cheung, and Lu 2022) 38.64 42.59 38.44 40.30
${}^{\star}$ (Jin, Lang, and Lei 2023) 41.50 40.80 31.40 39.40
@@
${}^{\star}$ (Xiang, Ding, and Han 2020) 39.30 39.60 24.20 36.20
RIDE (Wang et al. 2021b) (LC) 44.79 40.69 31.97 40.32
RIDE (Wang et al. 2021b) (CC) 44.38 40.59 32.99 40.35
${}^{\star}$ (Cui et al. 2023) 40.30 44.40 34.70 41.00
@@
  • Note: CC, cosine classifier. LC, linear classifier. The backbone is ResNet-152. Others are the same as Table 1.
Comparison Results: The results on CIFAR100-LT are shown in Table 1. H2T can further improve both single and multi-expert models. Applying H2T to the vanilla training of DR with CE loss (DR+H2T) enhances the performance by a significant margin, surpassing most recent methods. The biggest improvement exceeds 2%. H2T can also further boost the performance of two-stage SOTA methods, i.e., MisLAS and GCL, albeit with less significant improvements than for DR. H2T also bolsters the efficacy of multi-expert models. For example, RIDE and ResLT integrating with H2T outperform the original methods by 1.01% and 0.87% on CIFAR100-LT with $\rho=100$.
@@
The results on large-scale datasets are summarized in Tables 2, 3, and 4. The baseline (CE loss) achieves high accuracy on frequent classes while the performance on rare classes is unsatisfactory. Comparatively, all methods, except CMO and multi-expert models, exhibit substantial improvements in the medium and tail classes but at the expense of reduced accuracy in head classes. CMO leverages the background of head classes to augment tail classes without reducing the number of training samples in head classes, thus enhancing model performance for both head and tail classes. However, CMO+CE shows less competitive results in improving tail class performance. ABL also adjusts the decision boundary, but its effectiveness is not as pronounced as H2T. It is worth noting that all comparison methods necessitate training from scratch. H2T can outperform the comparison methods on most datasets by re-fining the classifier with only basic CE loss and several training epochs. For example, on iNaturalist 2018, DR+H2T achieves 71.81%, which is the highest among all other comparison methods. Furthermore, H2T can deliver consistent performance improvements for both single and multi-expert models, particularly for tail and medium classes. Nevertheless, the accuracy improvements of cosine classifier (CC) are less pronounced than those of linear classifier (LC). For example, we can observe that H2T improves GCL using CC much less than DR and MisLAS that use LC. One reason is that the soft margin in GCL allowing samples to fall in the margin between classes can alleviate the decision boundary bias compared with hard margins such as LDAM (Li et al. 2023). This phenomenon also confirms our motivation. Table 4 further compares the improvements of H2T on RIDE with LC (by 1.06%) and CC (0.95%). The margin on LC is more profound.
4.4 Further Analysis
This section visualizes the decision boundary of H2T on each class and investigates the effects of the sampler for the fusing branch and the fusion ratio $p$ mentioned in Section 3.3. All experiments are performed with DR+H2T.
[…]
Figure 5: Decision boundary Comparison of Class 0 and 8 without H2T v.s. with H2T (namely, DR and DR+H2T).
Figure 6: Change of accuracy on different splits w.r.t. $p$.
The Influence of Sampler for Fusing Branch: We compare different sampling strategies for the fusing branch, including class-balanced sampling (CS), reverse sampling (RS) that is tail-biased, and instance-wise sampling (IS) that is head-biased. The results are shown in Table 5. DR uses cRT to fine-tune the classifier with the class-balanced sampling data, which enlarges the medium and tail class expands by pushing the decision boundary in the direction of compressing head classes. Additionally, DR uses CE loss that does not consider class margin. These lead to a significant boost in medium and tail classes while reducing the classification accuracy on head classes. RS augments classes of all scales using head class samples with the lowest probability. CS makes all classes augmented with other classes in an equal probability. IS allows more features from the medium and tail classes to be fused with head features. All sampling strategies adjust the decision boundaries towards narrowing the head class distribution. RS has less damage to the performance of head classes while the improvement brought about by IS to the tail and medium classes is more prominent. As the classifier tuning stage cannot change the embedding distribution, the performance on medium and tail classes is further improved but with performance degradation of head classes.